Please remember to round your answer to two decimal places.

Problem 10-03 MIRR

A project has an initial cost of $43,425, expected net cash inflows of $12,000 per year for 10 years, and a cost of capital of 8%. What is the project's MIRR? Do not round intermediate calculations. Round your answer to two decimal places.______%

Problem 10-04 Profitability Index

A project has an initial cost of $61,000, expected net cash inflows of $12,000 per year for 11 years, and a cost of capital of 9%. What is the project's PI? Do not round your intermediate calculations. Round your answer to two decimal places. ________

Problem 10-05 Payback

A project has an initial cost of $59,050, expected net cash inflows of $13,000 per year for 8 years, and a cost of capital of 13%. What is the project's payback period? Round your answer to two decimal places._______ years

Problem 10-06 Discounted Payback

A project has an initial cost of $60,000, expected net cash inflows of $14,000 per year for 7 years, and a cost of capital of 13%. What is the project's discounted payback period? Round your answer to two decimal places.________ years

Please remember to round your answer to two decimal places.
Answer #1

Answer 10-03:

MIRR 14.88%

Workings:

Year Initial cost Operating cash flow Cash flows MIRR 0 1 2 4 6 7 9 10 ($43,425) $12,000$12,000 $12,000 $12,000 $12,000 $12,000 $12,000 $12,000 $12,000 $12,000 ($43,425)$12,000$12,000 $12,000 $12,000 $12,000 $12,000 $12,000 $12,000 $12,000 $12,000 14.88%

Answer 10-04:

Initial cash outflow = $61,000,

Expected net cash inflows per year for 11 years = $12,000

PV Factor for a One-Dollar Annuity Discounted at 9% for 11 Periods = [1 - 1 / (1 + k) n] / k

= (1 - 1/ (1+9%) 11) / 9% = 6.805191

PV of annual cash flows = $12,000 * 6.805191 = $81,662.29

PI = PV of annual cash flows / Initial cash outflow = $81,662.29 / $61,000 = 1.3387 = 1.34

Profitability Index 1.34

Answer 10-05:

Since project's annual cash flows are uniform:

Project's payback period = Initial cost / Annual cash flow

= $59,050 / $13,000 = 4.5423 = 4.54 years

Projects payback period 4.54 years

Answer 10-06:

Projects discounted payback period 6.68 years

Workings:

Year Initial cost Operating cash flow Cash flows PV factor [1/(1+13%)^year] Discounted cash flows Cumulative discounted cash flow Discounted payback period 0 2 4 6 7 ($60,000 14,000$14,000 $14,000 $14,000 $14,000$14,000 $14,000 ($60,000S14,000 S14,000S14,000 S14,000$14,000 S14,000 $14,000 1 0.884956 0.783147 0.693050 0.613319 0.542760 0.480319 0.425061 ($60,000) $12,389 $10,964 $9,703 $8,586 $7,599 $6,724$5,951 ($10,759)($4,034)$1,917 ($60,000) ($47,611) 6.68 Years (S36,647)($26,944)($18,357)